Self-Study Problem 5.03 Susan Wilson has $11,000 that she can deposit into a savings account for five years. Bank A compounds interest annually, Bank B twice a year, and Bank C quarterly Each bank has a stated interest rate of 4 percent. What account balance would Susan have at the end of the fifth year if she left all the interest paid on the deposit in each bank? (Round answers to 2 decimal places, e.g. 52.75.) Bank A Bank B Bank C Future Value
Answer #1

A:

We use the formula:
A=P(1+r/100)^n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$11000*(1.04)^5

=$11000*1.216652902

=$13383.18(Approx).

B:

We use the formula:
A=P(1+r/200)^2n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$11000*(1+0.04/2)^(2*5)

=$11000*1.21899442

=$13408.94(Approx).

C:

We use the formula:
A=P(1+r/400)^4n
where
A=future value
P=present value
r=rate of interest
n=time period.

A=$11000*(1+0.04/4)^(4*5)

=$11000*1.22019004

=$13422.09(Approx).